# GlobalMood: A cross-cultural benchmark for music emotion recognition

### v1

@@ 4.2.1 Cross-cultural ratings across the entire set
Having identified the top 20 mood terms for each language, we next gathered exhaustive ratings for all 1,180 songs. We recruited 1,741 new participants (see Section 3.1) who listened to the 15-second excerpts and rated how effectively each except conveyed a given mood term (1–5 scale). For each stimulus, participants evaluated seven randomly selected mood terms from the relevant language set. This systematic approach ensured that, on average, each song in each language received 8.38 (SD = 2.40) unique participant ratings, resulting in an extensive collection of 988,925 ratings spanning 1,180 songs across five languages.
@@ 4.2.2 Is ‘happy’ in my language the same ‘happy’ in your language?
To investigate differences in how each culture interprets these terms, we constructed 100 rating vectors (5 languages $\times$ 20 mood terms per language). Each vector was 1,180-dimensional, capturing the mean rating per term for each of the 1,180 songs. We then performed non-metric multidimensional scaling (MDS) using correlation as the distance metric, projecting these vectors in a two-dimensional space. In this mood ‘space,’ terms that position close to one another—even those from different languages—reflect similar rating patterns across the musical examples, suggesting comparable mood or emotional interpretations across cultures.
Figure 2A visualizes this mood space. The terms cluster into two main regions: high arousal and high valence (e.g., happy, energetic, and lively; upper region of the figure) and low arousal that span positive valence (e.g., peaceful; bottom left) to negative (e.g., sad; bottom right). Notably, many translated “equivalents” appear close together, which might suggest a general cross-cultural consensus.
However, examining six commonly shared terms that have direct translations in at least four of the five languages (fun, happy, rhythmic, love, sad, and calm) revealed varying degrees of cross-cultural agreement (see Figure 2B). For each of these terms, we computed mean correlations across language pairs and additionally reported the within-country measurement error ($r_{\text{within}}$), calculated using split-half reliability with Spearman-Brown formula (i.e., inter-country agreement across terms).
The term calm showed the highest average agreement ($r$ = 0.52 [0.49, 0.55]; $r_{\text{within}}$ = 0.49 [0.38, 0.57]), followed by fun ($r$ = 0.46 [0.39, 0.53]; $r_{\text{within}}$ = 0.44 [0.27, 0.54]), love ($r$ = 0.44 [0.41, 0.47]; $r_{\text{within}}$ = 0.47 [0.33, 0.66]), sad ($r$ = 0.41 [0.37, 0.45]; $r_{\text{within}}$ = 0.43 [0.32, 0.58]), rhythmic ($r$ = 0.38 [0.30, 0.45]; $r_{\text{within}}$ = 0.43 [0.25, 0.53]), and notably happy ($r$ = 0.37 [0.28, 0.45]; $r_{\text{within}}$ = 0.45 [0.24, 0.59]).
Overall, considering within-country agreement (mean $r_{\text{within}}$ = 0.43–0.48), most of these terms were comparable in their cross-country agreement (mean $r$ = 0.39–0.52). However, despite being considered a basic universal human emotion [34], happy, exhibited a considerable gap between cross- and within-country agreement. This emphasizes the necessity of incorporating diverse cultural perspectives when modeling nuanced musical mood responses. Reliance on either dictionary translation or LLM-based translation alone could overlook important, context-specific nuances in mood perception—particularly relevant when building models for global audiences.
@@
Recent benchmarks have evaluated the capabilities of audio LLMs across various downstream MIR tasks, but these evaluations have also been restricted to English [22]. Our cross-cultural dataset provides a unique opportunity to assess their performance across multiple languages. We evaluated both closed-source (Gemini) and open-source (CLAP) models against our GlobalMood multilingual human ratings (see Section 3.3 for model details).
@@
For Gemini (Figure 3A), we replicated our human study protocol by prompting the model with ‘Rate from a scale of 1 to 5 how well this song expresses or conveys the emotion […]’ in each of the five native languages. We then evaluated how well the model’s outputs predict human judgments. We observed consistent improvement with each new model version. The earliest model, 1.5 Flash, demonstrated modest alignment with human judgments, achieving an average cross-country correlation of $r$ = 0.34 (95% CI = [0.27, 0.42]). This model struggled particularly with Korean language ratings ($r$ = 0.27 [0.21, 0.33]). The subsequent 2.0 Flash version bridged this gap in Korean ($r$ = 0.42 [0.36, 0.47]), and achieved a substantially higher overall cross-country correlation of $r$ = 0.42 [0.36, 0.48]. The latest 2.5 Pro model demonstrated another leap with a cross-country average correlation of $r$ = 0.50 [0.45, 0.55]. Few-shot approach with 10 human-rated examples using 2.5 Pro did not improve the results ($r$ = 0.47 [0.42, 0.52]). This consistent upward trajectory across model iterations provides compelling evidence that these systems are progressively developing more sophisticated, human-like capabilities for understanding musical emotions across diverse linguistic and cultural contexts.
@@
This level of correlations between Gemini and human ratings are on par with algorithms specifically designed for estimating mood, such as Spotify’s high-level feature estimation [10, 35]. Moreover, given the subjective nature of mood perception in music (where even humans often disagree), the latest Gemini model already reaches human-level performance, matching the theoretical upper bound defined by inter-rater human agreement (gray dashed lines in Figure 3).
@@
For CLAP, an open-source alternative (Figure 3B), we evaluated both a zero-shot approach and a fine-tuned version trained on our GlobalMood dataset (see Section 3.3 for fine-tuning details). The zero-shot approach measured the similarity between the mood term’s text embedding and the song’s audio embedding. This zero-shot CLAP performed poorly (mean $r$ = 0.08 [0.03, 0.13]), while fine-tuning with GlobalMood substantially improved the performance (mean $r$ = 0.31 [0.19, 0.44]). As a control experiment, we also fine-tuned CLAP using a dataset where all non-English mood terms were first translated into English using an LLM without any musical context (instead of original mood terms collected from native speakers). This translation-based approach failed to improve performance (mean $r=0.13$ [-0.05, 0.32]), showing that the performance gains from fine-tuning come from culturally-specific information rather than from mere increase in data volume.
@@ 5 Discussion
We introduce GlobalMood, a novel cross-cultural benchmark dataset comprising mood descriptors and ratings collected from a large and globally diverse participant pool through a data-driven approach. Consistent with previous research in music perception [36, 37, 38, 39, 40] and music emotion [10, 41, 11, 12, 42], our findings highlight both cross-cultural similarities and differences. Specifically, we demonstrate that music mood descriptors across languages are broadly organized around clusters relating to valence and arousal [7, 34].
@@
Relying on dictionary translations for music mood terms may face limitation (e.g. [41]), where cross-lingual agreement varied across terms. For instance, despite happy being basic emotion, it exhibited relatively large gap between cross- and within country agreement. This highlights the need for incorporating multilingual mood descriptors from diverse annotators. Expanding the dataset to include additional languages [27, 43] and a broader range of musical styles, including works from different historical periods, will be critical for developing more robust tools that can generalize across diverse cultural contexts. Our selection of popular songs often includes lyrics and participants’ understanding of the lyrics may also have influenced their mood ratings. Future work could include instrumental music to better isolate mood judgments based solely on acoustic features.
@@
Notably, fine-tuning CLAP on cross-cultural data significantly boosted performance in non-English contexts, highlighting the value of GlobalMood. These findings point to the potential for culturally-sensitive MER systems, moving beyond one-size-fits-all models. Future work could enhance this further through prompting, more diverse data augmentation techniques, and exploring other LLM architectures.
@@
The tagging pipeline we adopt is quick to deploy across cultures and uses a bottom-up approach that helps minimize researcher bias. This offers an advantage over predefined mood lists, which often fail to capture the culturally grounded meanings of mood terms [44]. However, it has drawbacks: early participants can bias the tag pool, redundant or ambiguous tags require manual cleanup, and there’s no theoretical guarantee that all tags reflect clear mood concepts (e.g., étranger in French). Unlike structured systems for crowdsourcing tags (e.g., TagATune [45]), our approach trades control for flexibility and cultural responsiveness.
@@
Beyond MIR, the GlobalMood dataset and associated pipeline also offer timely implications for broader multimodal and cross-lingual research, particularly in Natural Language Processing (NLP) communities [46]. As NLP increasingly tackles multimodal tasks involving audio–text modeling and cross-cultural applications, our dataset may provide a useful benchmark for evaluating language–audio models across diverse contexts. The iterative annotation pipeline can also serve as an effective framework for collecting representative samples and annotations in other domains [14, 47, 13, 28].
@@ 6 Ethics Statement
We conducted our human experiment according to ethical best practices. All participants recruited via Prolific or CINT provided informed consent based on an approved protocol (Max Planck Ethics Council #202142). Participant data was collected anonymously (except for Prolific or CINT IDs), and all published data are fully anonymized. Models were accessed through commercial APIs, with fine-tuning performed locally. Code and training checkpoints will be provided in the published version. Conducting cross-cultural research requires sensitivity to diverse ethical considerations [48], and this study was designed accordingly. We acknowledge that machine learning models and training datasets may contain biases originating from participants, data selection processes, and the models themselves. Despite these potential biases, our study aims to address critical limitations and societal risks within current MIR technologies, which predominantly focus on English-language music and raters despite serving diverse global populations [49, 44, 50].

### v2

@@ 4.2.1 Cross-cultural ratings across the entire set
Having identified the top 20 terms for each language, we next gathered exhaustive ratings for the entire 1,180 songs of GlobalMood. We recruited 1,741 new participants (see Section 3.1) who listened to the 15-second excerpts and rated how effectively each excerpt conveyed a given emotion and mood term (1–5 scale). For each stimulus, participants evaluated seven randomly selected terms from the relevant language set. This systematic approach ensured that, on average, each song in each language received 8.38 (SD = 2.40) unique participant ratings, resulting in an extensive collection of 988,925 ratings spanning across five languages.
@@ 4.2.2 Is ‘happy’ in my language the same ‘happy’ in your language?
To investigate differences in how each culture interprets these terms, we constructed 100 rating vectors (5 languages $\times$ 20 terms per language). Each vector was 1,180-dimensional, capturing the mean rating per term across the 1,180 song set. We then performed non-metric multidimensional scaling (MDS) using correlation as the distance metric, projecting these vectors in a two-dimensional space. In this emotion ‘space,’ terms that position close to one another—even those from different languages—reflect similar rating patterns across the musical examples, suggesting comparable emotional interpretations across cultures.
Figure 2A visualizes this emotion space. The terms cluster into two main regions: one region of high arousal and high valence (e.g., happy, energetic, and lively; upper region of the figure) and a second region of low arousal that spans positive valence (e.g., peaceful; bottom left) to negative (e.g., sad; bottom right). Notably, many translated ‘equivalents’ appear close together, which might suggest a general cross-cultural consensus on what music evokes what emotions.
However, examining six commonly shared terms that have direct translations in at least four of the five languages (fun, happy, rhythmic, love, sad, and calm) revealed varying degrees of cross-cultural agreement (see Figure 2B). For each of these terms, between-country agreement ($r_{\text{between}}$) was computed as the average of pairwise correlation coefficients, while within-country agreement ($r_{\text{within}}$) was calculated using split-half reliability with Spearman-Brown formula. Effectively, $r_{\text{within}}$ serves as measurement error to compare as baselines when evaluating $r_{\text{between}}$.
The term calm showed the highest average agreement ($r_{\text{between}}$ = 0.52 [0.49, 0.55]; $r_{\text{within}}$ = 0.49 [0.38, 0.57]), followed by fun ($r_{\text{between}}$ = 0.46 [0.39, 0.53]; $r_{\text{within}}$ = 0.44 [0.27, 0.54]), love ($r_{\text{between}}$ = 0.44 [0.41, 0.47]; $r_{\text{within}}$ = 0.47 [0.33, 0.66]), sad ($r_{\text{between}}$ = 0.41 [0.37, 0.45]; $r_{\text{within}}$ = 0.43 [0.32, 0.58]), rhythmic ($r_{\text{between}}$ = 0.38 [0.30, 0.45]; $r_{\text{within}}$ = 0.43 [0.25, 0.53]), and notably happy ($r_{\text{between}}$ = 0.37 [0.28, 0.45]; $r_{\text{within}}$ = 0.45 [0.24, 0.59]).
Overall, considering within-country agreement (mean $r_{\text{within}}$ = 0.43–0.48), most of these terms were comparable in their between-country agreement (mean $r_{\text{between}}$ = 0.39–0.52). However, despite being considered a basic universal human emotion [37], happy exhibited a considerable gap between between- and within-country agreement. This emphasizes the necessity of incorporating diverse cultural perspectives when modeling nuanced musical emotional responses. Reliance on either dictionary translation or LLM-based translation alone could overlook important, context-specific nuances in emotion and mood perception—particularly relevant when building models for global audiences.
@@
Recent benchmarks have evaluated the capabilities of audio LLMs across various downstream MIR tasks, but these evaluations have also been restricted to English [25]. We evaluated both closed-source (Gemini) and open-source (CLAP) models against our GlobalMood multilingual human ratings (see Section 3.3 for model details).
@@
For Gemini (Figure 3A), we replicated our human study protocol by prompting the model with ‘Rate from a scale of 1 to 5 how well this song expresses or conveys the emotion […]’ independently for the five native languages. We then evaluated how well the model’s outputs predict human judgments. We observed consistent improvement with each new model version. The earliest model, 1.5 Flash, demonstrated modest alignment with human judgments (mean correlation across countries: $r$ = 0.34 (95% CI = [0.27, 0.42]). This model struggled particularly with Korean language ratings ($r$ = 0.27 [0.21, 0.33]). The subsequent 2.0 Flash version bridged this gap in Korean ($r$ = 0.42 [0.36, 0.47]), and achieved a substantially higher mean correlation of $r$ = 0.42 [0.36, 0.48]. The latest 2.5 Pro model demonstrated another leap with a mean correlation of $r$ = 0.50 [0.45, 0.55]. Few-shot approach with 10 human-rated examples using 2.5 Pro did not improve the results ($r$ = 0.47 [0.42, 0.52]). This consistent upward trajectory across model iterations provides compelling evidence that these general-purpose systems are progressively developing more sophisticated understanding of musical emotions across diverse linguistic and cultural contexts.
@@
This level of correlations between Gemini and human ratings are on par with algorithms specifically designed for MER, such as Spotify’s mood estimation [12, 38]. Moreover, given the subjective nature of emotion and mood perception in music (where even humans often disagree), the latest Gemini model already reaches human-level performance, matching the theoretical upper bound defined by inter-rater human agreement (gray dashed lines in Figure 3).
@@
For CLAP, an open-source alternative (Figure 3B), we evaluated both a zero-shot approach and a fine-tuned version trained on our GlobalMood dataset (see Section 3.3 for fine-tuning details). The zero-shot approach measured the similarity between the emotion term’s text embedding and the song’s audio embedding. This zero-shot CLAP performed poorly (mean $r$ = 0.08 [0.03, 0.13]), while fine-tuning with GlobalMood substantially improved the performance (mean $r$ = 0.31 [0.19, 0.44]). As a control experiment, we also fine-tuned CLAP using a dataset where all non-English emotion terms were first translated into English using an LLM without any musical context (instead of original emotion terms collected from native speakers). This translation-based approach failed to improve performance (mean $r=0.13$ [-0.05, 0.32]), showing that the performance gains from fine-tuning come from culturally-specific information rather than from mere increase in data volume.
@@ 5 Discussion
We introduce GlobalMood, a novel cross-cultural benchmark dataset comprising emotion and mood descriptors and ratings collected from a large and globally diverse participant pool through a data-driven approach. Consistent with previous research in music consumption and perception [39, 40, 41, 42, 43, 44] and music emotion [12, 45, 13, 14, 46], our findings highlight both cross-cultural similarities and differences. Specifically, we demonstrate that music emotion descriptors across languages are broadly organized around clusters relating to valence and arousal [8, 37].
@@
Relying on dictionary translations for music emotion terms may face limitations (e.g., [45]), where cross-lingual agreement varied across terms. For instance, despite happy being a basic emotion, it exhibited a relatively large gap between cross- and within-country agreement. This highlights the need for incorporating multilingual descriptors from diverse annotators. Our future work aims to expand GlobalMood to reach over 20 languages [30, 47] and a broader range of musical styles, including works from different historical periods, which will be critical for developing more robust tools that can generalize across diverse cultural contexts.
@@
Notably, fine-tuning CLAP on cross-cultural data significantly boosted performance in non-English contexts, highlighting the value of GlobalMood. These findings point to the potential for culturally-sensitive MER systems, moving beyond one-size-fits-all models. Future work should investigate more varied prompting and data augmentation techniques, and alternative LLM architectures.
@@
Our tagging pipeline is quick to deploy across cultures and uses a bottom-up approach that helps minimize researcher bias. This offers an advantage over predefined term lists, which often fail to capture the cultural subtleties of emotion terms [48]. However, it has drawbacks: early participants can bias the tag pool and there’s no theoretical guarantee that all tags reflect clear emotion concepts (e.g., ‘étranger’ in French). Unlike structured systems for crowdsourcing tags (e.g., TagATune [49]), our approach trades control for flexibility and cultural responsiveness.
@@
Beyond MIR, the GlobalMood and associated pipeline offer timely implications for broader multimodal and cross-lingual research, particularly in Natural Language Processing (NLP) communities [50]. As NLP increasingly tackles multimodal tasks involving audio–text modeling and cross-cultural applications, our dataset may provide a useful benchmark for evaluating language–audio models across diverse contexts. The iterative annotation pipeline can also serve as an effective framework for collecting representative samples and annotations in other domains [16, 51, 15, 31].
6 Acknowledgments
H.L. was funded by the Max Planck Society. M.P. was partially supported by the NYUAD Center for Interacting Urban Networks (CITIES), funded by Tamkeen under the NYUAD Research Institute Award CG001.
@@ 7 Ethics Statement
We conducted our human experiment according to ethical best practices. All participants recruited via Prolific or CINT provided informed consent based on an approved protocol (Max Planck Ethics Council #202142). Participant data was collected anonymously (except for Prolific or CINT IDs), and all published data are fully anonymized. Models were accessed through commercial APIs, with fine-tuning performed locally. Conducting cross-cultural research requires sensitivity to diverse ethical considerations [52], and this study was designed accordingly. We acknowledge that machine learning models and training datasets may contain biases originating from participants, data selection processes, and the models themselves. Despite these potential biases, our study aims to address critical limitations and societal risks within current MIR technologies, which predominantly focus on English-language music and raters despite serving diverse global populations [53, 48, 54].